# NILE [Uncaptioned image] : Internal Consistency Alignment in Large Language Models

### v1

@@ Abstract
As a crucial step to enhance LLMs alignment with human intentions, Instruction Fine-Tuning (IFT) has a high demand on dataset quality. However, existing IFT datasets often contain knowledge that is inconsistent with LLMs’ internal knowledge learned from the pre-training phase, which can greatly affect the efficacy of IFT. To address this issue, we introduce NILE (iNternal consIstency aLignmEnt) framework, aimed at optimizing IFT datasets to unlock LLMs’ capability further. NILE operates by eliciting target pre-trained LLM’s internal knowledge corresponding to instruction data. The internal knowledge is leveraged to revise the answer in IFT datasets. Additionally, we propose a novel Internal Consistency Filtering (ICF) method to filter training samples, ensuring its high consistency with LLM’s internal knowledge. Our experiments demonstrate that NILE-aligned IFT datasets sharply boost LLM performance across multiple LLM ability evaluation datasets, achieving up to $66.6\%$ gain on Arena-Hard and $68.5\%$ on Alpaca-Eval V2. Further analysis confirms that each component of the NILE framework contributes to these substantial performance improvements, and provides compelling evidence that dataset consistency with pre-trained internal knowledge is pivotal for maximizing LLM potential.
@@
Minda Hu1, Qiyuan Zhang2, Yufei Wang3, Bowei He2, Hongru Wang1, Jingyan Zhou1 Liangyou Li3, Yasheng Wang3, Chen Ma2, Irwin King1 1The Chinese University of Hong Kong 2City University of Hong Kong   3Huawei Noah’s Ark Lab {mindahu21, king}@cse.cuhk.edu.hk
@@ 1 Introduction
Instruction Fine-Tuning (IFT), which fine-tunes LLMs on instruction-response pairs, has been proven to be an effective and crucial method to enhance the capabilities and controllability of large language models (LLMs) (Touvron et al., 2023; Dubey et al., 2024; Zhang et al., 2023). Most IFT approaches predominantly focus on the quantity and diversity of datasets, based on the assumption that greater size of instruction-response pairs would lead to better performance (Honovich et al., 2023; Wang et al., 2023a; Taori et al., ; Chiang et al., 2023; Sun et al., 2024). These approaches narrowly focus on data quantity while overlooking IFT’s core purpose: unlocking the latent capabilities of pre-trained LLMs. They fail to examine the crucial relationship between SFT training data and the underlying pre-trained models.
@@
Recent findings delve into this problem deeper, revealing that the alignment between underlying world knowledge derived from IFT datasets and the internal knowledge stored within the parameters of LLMs is crucial for determining their performance Kang et al. (2024). This alignment creates meaningful connections between IFT datasets and the downstream performance of specific LLMs, offering theoretical guidelines for the generation and selection of IFT datasets tailored to enhance the effectiveness of particular LLMs. Specifically, when world knowledge from IFT datasets conflicts entirely with LLM’s internal knowledge, these encoded associations are largely neglected, and the LLM tends to merely predict an intelligent “blind guess” by mimicking surface-level patterns such as text styles without truly understanding task intentions. Nevertheless, further experiments Ren et al. (2024) observe that only incorporating a certain degree of world knowledge that is incompatible with LLMs’ internal parameter knowledge into IFT datasets brings more performance boost. There are recent attempts (Li et al., 2024b; Cao et al., ; Li et al., 2024d; Liu et al., 2024; Li et al., 2024a) to maintain consistent formatting at a superficial level such as text styles, but they still overlook the need for managing internal consistency.
Inspired by previous findings, we propose a novel framework called NILE[Uncaptioned image] (INternal ConsIstency aLignmEnt). To tackle the aforementioned issues, NILE is designed to generate and select better IFT datasets considering the consistency between internal parameter knowledge in LLMs and world knowledge in IFT datasets. Specifically, we leverage the off-the-shelf LLMs to revise given IFT datasets based on the internal parameter knowledge and then select the samples considering the relationship between the implicit world knowledge and internal parameter knowledge. It is important to highlight that our method does not rely on any additional forms of supervision (i.e., human experts). To conclude, our contributions can be summarized as follows:
  • We propose NILE, a novel framework to generate and select better IFT datasets considering the consistency between internal parameter knowledge in LLMs and world knowledge in IFT datasets, as illustrated in Figure 1.
  • Through comprehensive ablation studies and empirical analysis, we demonstrate that maintaining consistency between IFT datasets and LLMs’ internal knowledge is crucial for unlocking model capabilities. Our results provide strong evidence that each component of NILE contributes to performance gains.
@@ 2.1 Data Synthesis in Instruction Tuning
Earlier research on instruction tuning has primarily focused on developing large, high-quality datasets curated by human experts (Wei et al., 2022; Wang et al., 2022). However, this process is often time-consuming and labor-intensive. Thus, several studies have explored the use of more advanced models or self-critique prompting methods (Wang et al., 2024; Zhang et al., 2024) to automatically generate instruction-tuning datasets. For examples, Self-Instruct (Wang et al., 2023a) leverage GPT-3 to expand asks to many diverse domains in the in-context learning manner while several recent studies directly use latest SOTA model to generate the response or reflect on current samples Mukherjee et al. (2023), such as WizardLM Xu et al. (2023) and Reflection-tuning Li et al. (2024a). In addition to focusing on quality side, another area of work aims to create more diverse and larger instruction-tuning datasets. For example, UltraChat Ding et al. (2023) defines specific scopes and systematically generates a wide range of instructions within each area. In contrast, Magpie (Xu et al., 2024) only feed the left-side templates up to the position reserved for user messages as input to generate more diverse user queries.
@@ 2.2 Data Selection in Instruction Tuning
Data selection (or revision) have been widely studied in large language model instruction tuning, considering the importance of data quality in model training (Li et al., 2024b; Cao et al., ; Li et al., 2024d; Zhou et al., 2024; Liu et al., 2024; Li et al., 2024a). Most previous studies fall into two categories: 1) relying on more powerful models or human experts to select better data (Zhou et al., 2024; Liu et al., 2024); 2) calculating the ppl gains considering generated samples and original samples (Li et al., 2024a). While both methods improve downstream performance, they face significant limitations, such as the high cost of human labeling. More importantly, such studies (Chen et al., ; Li et al., ; Sun et al., 2024) can not provide fundamental explanations regarding the key factors that define better instruction-tuning datasets. In contrast to these approaches, our work aligns the internal knowledge of LLMs with external world knowledge derived from ITF datasets, resulting in improved datasets that offer better explainability and transparency.
Figure 2: Overview of our NILE framework.
@@ 3 Method
Figure  2 demonstrates our framework NILE for increasing knowledge affinity between LLMs’ internal knowledge and instruction-tuning datasets. It can be divided into three parts: (1) Internal Knowledge Extraction (IKE), (2) Knowledge-aware Sample Revision (KSR), and (3) Internal Consistency Filtering (ICF). IKE accesses the memory of pretrained LLMs to sample their internal knowledge. KSR revises existing dataset samples by automatically infusing the sampled internal knowledge. ICF introduces a novel internal consistency measurement to filter out low-quality revisions from the second phase. Notably, our framework is decoupled from and easily integrated into current LLM alignment pipelines. In the following subsections, we introduce the above three components in detail. Implementation details of IKE, KSR, and ICF are listed in Appendix A.1.
@@ 3.1 Internal Knowledge Extraction
This step aims to effectively sample the internal knowledge from the target pre-trained LLM $\mathcal{M}$ for instructions in the original IFT dataset. Specifically, for original instruction tuning datasets, $\mathcal{D}^{o}$ has $n$ data samples formatted as $\mathcal{D}^{o}=\{q^{o}_{1...n},\ \mathrm{a}^{o}_{1...n}\}$. Here we define the concatenated sequence $q^{o}=\{\mathrm{instruction}^{o},\mathrm{input}^{o}\}$ as a complete instruction, and $a^{o}=\mathrm{answer}^{o}$ as its corresponding answer. We aim to guide $\mathcal{M}$ to output its corresponding internal knowledge $ik_{q^{o}}^{\mathcal{M}}$ related to $q^{o}$ by utilizing in-context learning techniques, i.e., few-shot prompting. Instead of fixed examples in the prompt, we first select relevant samples as examples, then prompt $\mathcal{M}$ to generate knowledge w.r.t. $q^{o}$.
Generate a list of related knowledge about the following Instruction and Input up to 500 words. Do not directly output the answer, but focus on the related knowledge required for answering the Input. Instruction: "$\mathbf{\{\mathrm{\textbf{instruction}}^{d}_{i}\}}$" Input: "$\mathbf{\{\mathrm{\textbf{input}}^{d}_{i}\}}$"
Table 1: Prompt for generating internal knowledge demonstration $ik^{d}_{i}$ related to $q^{d}_{i}$.
As shown in Figure 2, the process of IKE involves the following steps:
@@
    We first construct database index $\mathcal{F}^{demo}=\{(q^{d}_{1},ik^{d}_{1}),\ldots,(q^{d}_{m},ik^{d}_{m})\}$, which is a fix set of demonstrations contains $m$ samples. $\{q_{i}^{d}\}$ is randomly sampled from an IFT dataset. We then prompt a strong LLM as the teacher to generate corresponding knowledge $ik^{d}_{i}$ for each $q^{d}_{i}$ as shown in Table 1. Details are provided in Appendix A.1.3.
@@
    For each $q^{o}$, we select $k$ few-shot examples $f^{\mathcal{R}}(q^{o})=\{(q^{\mathcal{R}}_{1},ik^{\mathcal{R}}_{1}),\ldots,(q^{\mathcal{R}}_{k},ik^{\mathcal{R}}_{k})\}$ from $\mathcal{F}^{demo}$ by the semantic similarity between $q^{o}$ and $\{q_{i}^{d}\}$. We employ information retrieval algorithms, such as BM25, as the retriever $\mathcal{R}$ to compute the semantic similarity.
@@
    The retrieved demonstration samples $f^{\mathcal{R}}(q^{o})$ are fed into the target LLM $\mathcal{M}$ to enable in-context learning. This allows the LLM to accurately expose its internal knowledge $ik^{\mathcal{M}}_{q}$ related to the original instruction $q^{o}$ by utilizing the prompt design in Table 2.
{few_shot_prompt $\mathbf{(q^{\mathcal{R}}_{1...k},ik^{\mathcal{R}}_{1...k})}$ Instruction: $\mathbf{\{\mathrm{\textbf{instruction}}^{o},\mathrm{\textbf{input}}^{o}\}}$ Related Knowledge:
@@ 3.2 Knowledge-aware Sample Revision
After obtaining an accurate sampling $ik^{\mathcal{M}}_{q^{o}}$ of the target LLM’s internal knowledge for each original instruction $q^{o}$, we design a prompt for the revisor LLM agent $\mathcal{A}_{r}$ to infuse $ik^{\mathcal{M}}_{q^{o}}$ into the current instruction and get the revised answer $a^{ik}$. The prompt for KSR is displayed in Table 3.
Provide a better response based on "$\mathbf{\{a^{o}\}}$" to comply with given instruction, input, and related knowledge. Instruction: $\mathbf{\{\mathrm{\textbf{instruction}}^{o}\}}$ Input:$\mathbf{\{\mathrm{\textbf{input}}^{o}\}}$ Related Knowledge: $\{ik^{\mathcal{M}}_{q^{o}}\}$ Please directly output the improved response.
@@
This step aims to enhance affinity between the target model $\mathcal{M}$’s internal knowledge $ik^{\mathcal{M}}_{q^{o}}$ and the original answer $a^{o}$ from $\mathcal{D}^{o}$ with world knowledge, resulting an improved answer $a^{ik}$.
@@ 3.3 Internal Consistency Filtering
In this stage, we evaluate the effectiveness of KSR by comparing the quality of the revised answer $a^{ik}$ with the original answer $a^{o}$. Drawing inspiration from IFD and PMI Li et al. (2023), we introduce a novel metric called Internal Consistency Index (ICI) to quantify how well one answer promotes knowledge associations in the pretrained LLM $\mathcal{M}$.
@@
$\operatorname{ICI}_{\mathcal{M}}(q,a^{ik})=\frac{P_{\mathcal{M}}(a^{ik}\mid q,ik^{\mathcal{M}}_{q})}{P_{\mathcal{M}}(a^{ik}\mid q)},$ (2)
where $P_{\mathcal{M}}(a^{ik}\mid q)$ measures the associations between revised responses $a^{ik}$ and instructions $q$ alone, while $P_{\mathcal{M}}(a^{ik}\mid q,ik^{\mathcal{M}}_{q})$ captures the overall association strength between $a^{ik}$ and the combination of $q$ and its corresponding extracted internal knowledge $ik^{\mathcal{M}}_{q}$. To isolate the influence of $ik^{\mathcal{M}}_{q}$ on the revised answer $a$, we minimize the influence of $q$ in the ICI formulation by dividing $P_{\mathcal{M}}(a\mid q,ik^{\mathcal{M}}_{q})$ with $P_{\mathcal{M}}(a\mid q)$.
For samples with higher ICI values, the model more effectively integrates and leverages the explicitly provided internal knowledge when generating the revised answer, suggesting a stronger alignment between the revised answer and the model’s internal knowledge. Conversely, for samples with lower ICI values, providing internal knowledge may not benefit or could even hinder the generation of the revised answer, indicating that the revised answer does not have a strong association with what the model has learned internally, as suggested by Ren et al. (2024). Therefore, we employ a filtering mechanism ICF to filter out these redundant low ICI samples to an aligned dataset $\mathcal{D}^{aligned}$ for fine-tuning an aligned LLM $\mathcal{M}^{a}$ from $\mathcal{M}$. To control other variables in the experiment and ensure stable improvement, we revert to the original samples $(q,a^{o})$ when the ICI values of $(q,a^{ik})$ are lower than threshold $\beta$:

### v2

@@ Abstract
Recent advances show that the world knowledge in the Instruction Fine-Tuning (IFT) dataset, which is incompatible with LLMs’ internal knowledge, can greatly hurt the IFT performance. However, the effective integration and balancing of the internal knowledge of LLMs, acquired during pre-training, with existing IFT datasets remains a largely underexplored area of research. To address this gap, this work introduces NILE, a novel framework to optimize the effectiveness of IFT by adjusting IFT datasets through carefully aligning the world and internal knowledge. NILE employs a three-stage pipeline to effectively quantify and adjust consistency with the internal knowledge of target LLMs. Our analysis provides compelling evidence that balancing such consistency with pre-trained internal knowledge is pivotal for unleashing LLM potential, and confirms that NILE can systematically contribute to these substantial performance improvements. Experimental results demonstrate that NILE-aligned IFT datasets sharply boost LLM performance across multiple LLM ability evaluation datasets, achieving up to $66.6\%$ gain on Arena-Hard and $68.5\%$ on Alpaca-Eval V2.
@@
Minda Hu1, Qiyuan Zhang2, Yufei Wang3, Bowei He2, Hongru Wang4 Jingyan Zhou1, Liangyou Li3, Yasheng Wang3, Chen Ma2, Irwin King1 1The Chinese University of Hong Kong 2City University of Hong Kong 3Huawei Noah’s Ark Lab 4University of Edinburgh {mindahu21, king}@cse.cuhk.edu.hk
@@ 1 Introduction
Instruction Fine-Tuning (IFT), which fine-tunes Large Language Models (LLMs) on instruction-response pairs, has been proven to be an effective and crucial method to enhance the capabilities and controllability of LLMs (Touvron et al., 2023; Dubey et al., 2024; Zhang et al., 2023; Chen et al., 2024; Wang et al., 2025). Most IFT approaches predominantly focus on the quantity and diversity of datasets, based on the assumption that a greater size of instruction-response pairs would lead to better performance (Honovich et al., 2023; Wang et al., 2023b; Taori et al., ; Chiang et al., 2023; Sun et al., 2024). These approaches narrowly emphasize data quantity while overlooking IFT’s core purpose: unlocking the latent capabilities of pre-trained LLMs. They do not adequately consider underlying correlations between IFT datasets and LLMs, which is crucial to the efficacy of IFT Ren et al. (2024).
@@
A key factor influencing IFT performance is the level of internal consistency, i.e., the consistency between the world knowledge in IFT datasets and the internal knowledge embedded within LLM parameters Ren et al. (2024). When trained on totally unfamiliar data, i.e., data with low internal consistency, LLMs may only capture superficial correlations in instruction-response pairs, such as text styles, and tend to make “blind guesses” when faced with new queries Kang et al. (2024). Nonetheless, Ren et al. (2024) shows that merely maximizing internal consistency does not necessarily lead to optimal IFT performance. These works suggest that examining and curating the internal consistency of IFT datasets for the target pre-trained LLM is a promising direction for effective training. However, how to revise and balance the internal consistency level remains under-explored.
In this work, we propose a novel framework NILE[Uncaptioned image] (INternal ConsIstency aLignmEnt). NILE bridges the aforementioned research gap by flexibly improving existing IFT datasets in terms of internal consistency for the target pretrained LLM. Specifically, NILE addresses the problem through the following three steps: 1) Internal Knowledge Extraction. As a prerequisite, accurately extracting internal knowledge is crucial. We adopt in-context learning techniques with high-quality customized examples. 2) Knowledge-aware Sample Revision. To fully utilize existing data, we designed a revision step to improve the existing data with LLMs’ internal knowledge, resulting in a data sample with higher consistency. 3) Internal Consistency Filtering. Lastly, we developed a novel metric to measure the consistency level between the data sample and the LLM. By doing so, we can flexibly adjust the level of internal consistency of existing IFT data with any target pre-trained LLM to achieve optimized IFT performance.
It is important to highlight that our method does not rely on any additional forms of supervision (i.e., human experts). To conclude, our contributions can be summarized as follows:
  • We propose NILE, a novel framework to adjust and select better IFT datasets considering the consistency between internal parameter knowledge in LLMs and world knowledge in IFT datasets, as shown in Figure 1. To the best of our knowledge, we are among the first to leverage the concept of internal consistency for IFT data selection and generation. [^1] Corresponding NILE-revised IFT datasets can be found in https://huggingface.co/datasets/mindahu/NILE-IFT-Dataset.
  • Through comprehensive ablation studies and empirical analysis, we demonstrate that balancing consistency between IFT datasets and LLMs’ internal knowledge is crucial for unlocking model capabilities. Our results provide strong evidence that each component of NILE contributes to performance gains.
@@ 2.1 Data Synthesis in Instruction Tuning
Earlier research on instruction tuning has primarily focused on developing large, high-quality datasets curated by human experts (Wei et al., 2022; Wang et al., 2022). However, this process is often time-consuming and labor-intensive. Thus, several studies have explored the use of more advanced models or self-critique prompting methods (Wang et al., 2023a, 2024; Zhang et al., 2024; Pi et al., 2024b, a) to generate instruction-tuning datasets automatically. For example, Self-Instruct (Wang et al., 2023b) leverages GPT-3 to expand asks to many diverse domains in an in-context learning manner while several recent studies directly use the latest SOTA model to generate the response or reflect on current samples Mukherjee et al. (2023), such as WizardLM Xu et al. (2023) and Reflection-tuning Li et al. (2024a). In addition to focusing on the quality side, another area of work aims to create more diverse and larger instruction-tuning datasets. For example, UltraChat Ding et al. (2023) defines specific scopes and systematically generates a wide range of instructions within each area. In contrast, Magpie (Xu et al., 2024) only feeds the left-side templates up to the position reserved for user messages as input to generate more diverse user queries.
For complex reasoning tasks such as coding and mathematics, many efforts have been made to integrate human priors into data synthesis Zhou et al. (2025). KPDDS Huang et al. (2025) leverages key points and exemplar practices to synthesize mathematical reasoning-focused IFT datasets. Additionally, Case2Code Shao et al. (2024) introduces observations of input-output examples and program behaviors to infer underlying code implementations.
@@ 2.2 Data Selection in Instruction Tuning
Data selection (or revision) has been widely studied in large language model instruction tuning, considering the importance of data quality in model training (Li et al., 2024b; Cao et al., 2023; Li et al., 2024d; Zhou et al., 2024; Liu et al., 2024; Li et al., 2024a). Most previous studies fall into two categories: 1) relying on more powerful models or human experts to select better data (Zhou et al., 2024; Liu et al., 2024); 2) calculating the perplexity gains considering generated samples and original samples (Li et al., 2024a). While both methods improve downstream performance, they face significant limitations, such as the high cost of human labeling. More importantly, such studies (Chen et al., ; Li et al., ; Sun et al., 2024) can not provide fundamental explanations regarding the key factors that define better instruction-tuning datasets. In contrast to these approaches, our work aligns the internal knowledge of LLMs with external world knowledge derived from IFT datasets, resulting in improved datasets that offer better explainability and transparency.
Figure 2: Overview of our NILE framework. NILE consists of three main steps: Internal Knowledge Extraction (IKE), Knowledge-aware Sample Revision (KSR), and Internal Consistency Filtering (ICF).
@@ 3 Method
Figure  2 demonstrates our framework NILE for increasing knowledge affinity between LLMs’ internal knowledge and instruction-tuning datasets. It can be divided into three parts: (1) Internal Knowledge Extraction (IKE), (2) Knowledge-aware Sample Revision (KSR), and (3) Internal Consistency Filtering (ICF). The core contribution of our framework lies in our deliberate focus on internal consistency, which enables the process to function effectively. IKE accesses the memory of pretrained LLMs to sample their internal knowledge. KSR revises existing dataset samples by automatically infusing the sampled internal knowledge. ICF introduces a novel internal consistency measurement to filter out low-quality revisions from the second phase. In the following subsections, we introduce the above three components in detail. Implementation details of IKE, KSR, and ICF are listed in Appendix A.1.
@@ 3.1 Internal Knowledge Extraction
This stage aims to effectively sample the internal knowledge from the target pre-trained LLM $\mathcal{M}$ for instructions in the original IFT dataset $\mathcal{D}^{o}=\{(q^{o}_{i},\ a^{o}_{i})\}_{i=1}^{n}$, where $q^{o}_{i}$ is the concatenated query sequence of $\mathrm{\textbf{instruction}}^{o}_{i}$ and $\mathrm{\textbf{input}}^{o}_{i}$, and $a^{o}_{i}$ is the answer. Formally, we aim to sample the internal knowledge $ik_{i}$ corresponding to $q^{o}_{i}$ from $\mathcal{M}$ through in-context learning. Instead of using a fixed set of examples, we use following three-step strategy to provide the most relevant examples to better exert the internal knowledge from $\mathcal{M}$.
@@
    Demonstration set construction: we first randomly sample a subset of queries $\{q_{j}^{d}\}_{j=1}^{m}$ from an IFT dataset. Then, as shown in Table 1, a strong LLM (GPT-4 utilized in the experiments) is prompted to generate the corresponding knowledge snippet $ik^{d}_{j}$ for each $q_{j}^{d}$, resulting in a demonstration database index $\mathcal{F}^{demo}=\{(q^{d}_{j},ik^{d}_{j})\}_{j=1}^{m}$. Details are provided in Appendix A.1.3.
@@
    Example selection: for each query $q^{o}_{i}\in\mathcal{D}^{o}$, we select $k$ few-shot examples $f^{\mathcal{R}}(q_{i}^{o})=\{(q^{d}_{i_{t}},ik^{d}_{i_{t}})\}_{t=1}^{k}$ from $\mathcal{F}^{demo}$, where $({i_{t}})_{t=1}^{k}$ denotes the indices of top-$k$ example pairs ranked by the query semantic similarity between $\{q_{j}^{d}\}_{j=1}^{m}$ and $q^{o}_{i}$ from retriever $\mathcal{R}$. $\mathcal{R}$ is implemented by information retrieval algorithms such as BM25.
@@
    Internal knowledge generation: we formulate the prompt shown in Table 2 to $\mathcal{M}$ with few-shot examples $f^{\mathcal{R}}(q^{o}_{i})$ and the original instruction $q^{o}_{i}$. By this means, it can effectively exert internal knowledge $ik_{i}$ from the target LLM $\mathcal{M}$.
Generate a list of related knowledge about the following Instruction and input up to 500 words. Do not directly output the answer, but focus on the related knowledge required for answering the Input. Instruction: $\mathbf{\{\mathrm{\textbf{instruction}}^{d}_{j}\}}$ Input: $\mathbf{\{\mathrm{\textbf{input}}^{d}_{j}\}}$
Table 1: Prompt for demonstration set construction.
$\mathbf{\{(q^{d}_{i_{t}},ik^{d}_{i_{t}})\}_{t=1}^{k}}$ Instruction: $\mathbf{\{\mathrm{\textbf{instruction}}^{o}_{i},\mathrm{\textbf{input}}^{o}_{i}\}}$ Related Knowledge:
@@ 3.2 Knowledge-aware Sample Revision
After obtaining a relatively accurate sampling $ik_{i}$ of the target LLM’s internal knowledge (analyzed in Section 4.8), for each original instruction $q^{o}_{i}$, we design a prompt for the revisor LLM agent $\mathcal{A}_{r}$ to infuse $ik_{i}$ into the current instruction and get the revised answer $a^{ik}_{i}$. The prompt for KSR is displayed in Table 3.
[…]
@@
This step aims to enhance affinity between the target model $\mathcal{M}$’s internal knowledge $ik_{i}$ and the original answer $a^{o}_{i}$ from $\mathcal{D}^{o}$ with world knowledge, resulting an improved answer $a^{ik}_{i}$.
@@ 3.3 Internal Consistency Filtering
In this stage, we evaluate the effectiveness of KSR by comparing the quality of the revised answer $a^{ik}_{i}$ with the original answer $a^{o}_{i}$. Drawing inspiration from IFD and PMI Li et al. (2023), we introduce a novel metric called Internal Consistency Index (ICI) to quantify how well one answer promotes knowledge associations in the pretrained LLM $\mathcal{M}$.
@@
$\operatorname{ICI}_{\mathcal{M}}(q,a^{ik})=\frac{P_{\mathcal{M}}(a^{ik}\mid q,ik)}{P_{\mathcal{M}}(a^{ik}\mid q)},$ (2)
where $P_{\mathcal{M}}(a^{ik}\mid q)$ measures the associations between revised responses $a^{ik}$ and instructions $q$ alone, while $P_{\mathcal{M}}(a^{ik}\mid q,ik)$ captures the overall association strength between $a^{ik}$ and the combination of $q$ and its corresponding extracted internal knowledge $ik$. To isolate the influence of $ik$ on the revised answer $a^{ik}$, we minimize the influence of $q$ in the ICI formulation by dividing $P_{\mathcal{M}}(a^{ik}\mid q,ik)$ with $P_{\mathcal{M}}(a^{ik}\mid q)$.
For samples with higher ICI, the model more effectively integrates and leverages the explicitly provided internal knowledge when generating the revised answer, suggesting a stronger alignment between the revised answer and the model’s internal knowledge. Conversely, for samples with lower ICI, providing internal knowledge may not benefit or could even hinder the generation of the revised answer, indicating that the revised answer does not have a strong association with what the model has learned internally, as suggested by Ren et al. (2024). Therefore, we employ a filtering mechanism ICF to filter out these redundant low ICI samples to an aligned dataset $\mathcal{D}^{aligned}$ for fine-tuning an aligned LLM $\mathcal{M}^{a}$ from $\mathcal{M}$. To control dataset size in the experiment and ensure stable improvement, we revert to the original samples $(q,a^{o})$ when the ICI values of $(q,a^{ik})$ are lower than the threshold $\beta$:
@@
Here we use $\beta$ to control the degree of internal consistency in ICF.